OpenStudy (earthcitizen):

The barometer of a mountain hiker reads 930 mbars at the beginning of a hiking trip and 780 mbars at the end. Neglecting the effect of altitude on local gravitational acceleration, determine the verticl distance climed. Assume an average air density of 1.20kg/m^3

OpenStudy (earthcitizen):

\[P _{2}-P _{1}=\rho g h\]

OpenStudy (earthcitizen):

\[P _{2 }=930 mbars =93000 Pa, P _{1}= 780mbars=78000 Pa\]

OpenStudy (earthcitizen):

\[\rho = 1.20kg/m^3, g=9.81ms ^{-1} and h= ?\]

OpenStudy (earthcitizen):

\[\therefore, 93000-78000= (1.20)*(9.81)*(h) \]

OpenStudy (earthcitizen):

h= 1274m

OpenStudy (anonymous):

Why did you answer your own question? ;)

OpenStudy (anonymous):

can anyone answer , use the ideal gas law to explain the changes in temperature, pressure and volume of the coolant gas at the compressors stage of its refrigeration